3. A stone is dropped in a well. (a) The sound of the splash is heat t seconds later. If d represents the depth of the well find an expression for d. (b) Find d when t 3

I would like some help with this question if you can! Thank you!
Answer #1

2 (time to hit the water)

to Usound (time sound takes to travel the distance(depth))

t=t_1+t_2

2d d sound

2d Usoundt- dUsound

ondd2dvsoundt 2duund

v_{sound}^2gt^2+gd^2-2gdv_{sound}t=2dv_{sound}^2

v_{sound}^2gt^2+gd^2-2gdv_{sound}t-2dv_{sound}^2=0

d^2-left (2v_{sound}t+rac {2v_{sound}^2}{g} ight )d+v_{sound}^2t^2=0

soundt +nd